# Unraveling hydrodynamization using ultracold 1D gases

### v2

@@ II.1 Experiments and models
We carry out our theoretical analyses in the strong coupling Tonks–Girardeau (TG) limit ($\gamma\to\infty$) of Hamiltonian (1), where out-of-equilibrium momentum and rapidity distributions can be readily calculated following Bragg scattering quenches. As shown in Ref. [5], the behavior in that limit is qualitatively similar to the experimental observations of hydrodynamization and local prethermalization at finite $\gamma$. In our numerical calculations, we use the low-density regime of the lattice hard-core boson Hamiltonian,
${\cal H}_{\rm HCB}=-J\sum_{j=1}^{N_{s}-1}\big{(}\hat{b}^{\dagger}_{j+1}\hat{b}_{j}+{\rm H.c.}\big{)}+\sum_{j=1}^{L}U(z_{j})\hat{b}^{\dagger}_{j}\hat{b}_{j}\,,$ (3)
where $J$ is the hopping amplitude of a hard-core boson, and $N_{s}$ is the total number of lattice sites. The boson creation (annihilation) operator $\hat{b}^{\dagger}_{i}$ ($\hat{b}_{i}$) creates (annihilates) a boson at lattice site $i$, and the hard-core constraint is enforced by $\hat{b}_{i}^{2}=\hat{b}^{\dagger 2}_{i}=0$. The position of site $j$ on the lattice is set to be $z_{j}=(j-N_{s}/2)a$, with $a$ the lattice spacing. In the limit in which the site occupation $n_{j}=\langle\hat{b}^{\dagger}_{j}\hat{b}_{j}\rangle\to 0$ on all sites, the lattice hard-core boson Hamiltonian (3) is equivalent to the continuum Hamiltonian (1) in the TG limit [19]. The parameters of these two Hamiltonians are connected by $J=\hbar^{2}/(2ma^{2})$.
The main experimental observable is the momentum distribution $f(p)$, which is measured via a time-of-flight expansion [20] after the 1D interactions have been essentially turned off [3]. For the TG theory, $f(p)$ is written as (up to a normalization constant)
@@
II.2 Single Bragg pulse
To create out-of-equilibrium states that feature hydrodynamization, we apply a Bragg pulse quench to the system [5] with a $\lambda$ wavelength standing wave. That is, we turn on an additional potential,
@@ II.2 Single Bragg pulse
with wavenumber $k$ and amplitude $U_{\rm pulse}$, for a short period of time $t_{\rm pulse}$. In our experiments $k\approx k_{0}$, where $k_{0}$ is the wavenumber of the 2D lattice. After the pulse, we let the system evolve in the initial trapping potential [$U(z)$ in Eq. (2)] for different times $t_{\rm ev}$ and then measure the momentum distribution $f(p)$.
The effect of the Bragg pulse can be straightforwardly understood in the Raman-Nath regime ($t_{\rm pulse}\to 0$) for the homogeneous TG gas. The pulse in this regime, known as a Kapitza-Dirac pulse, “kicks” the particles and quasiparticles before they have had time to move. Such a kick generates a state in which the quasiparticles are in a superposition of plane waves that differ by even integer multiples of the Bragg momentum. The many-quasiparticle (noninteracting fermion) wave function changes from that of a Fermi sea $|\Psi_{0}\rangle=\prod_{|\theta|<\theta_{F}}\hat{c}^{\dagger}_{\theta}|0\rangle$ to
@@
$I_{n}(-ib)$ are the modified Bessel functions of the first kind, where the integer $n$ is the order of the Bragg peak and $b$ depends on the area of the Bragg pulse [23]. In a homogeneous system, the rapidity distribution $f(\theta)$ is conserved after the quench, but the momentum distribution $f(p)$ for the bare particles evolves in time. The time evolution of the relative phases of different $\hat{c}^{\dagger}_{\theta}$ enters $f(p)$ via the Bose-Fermi mapping, so $f(p)$ changes in time despite $f(\theta)$ being constant. That is, although there is no redistribution of energy among the quasiparticles, there is a redistribution of energy among different momentum modes of the particles.
The largest relative energy difference between significantly populated quasiparticle states sets the fastest rate that $f(p)$ can evolve. This is the “hydrodynamization coherence” time $T^{c}_{\rm hd}$. Specifically, if the $n=0$ and ±$\pm 1$ Bragg peaks are the mostly populated ones, $\Delta E\simeq(2\hbar k)^{2}/2m$ and we have the characteristic frequency $\omega^{c}_{\rm hd}=\hbar(2k)^{2}/2m$ and corresponding period
$T^{c}_{\rm hd}=\frac{2\pi}{\omega^{c}_{\rm hd}}=\frac{\pi m}{\hbar k^{2}}.$ (9)
Figure 1: Experimental results from Ref. [5] for $\bar{\gamma}_{0}=2.3$ after a $6\,\mu$s Bragg pulse. (a) Time evolution of $E_{\bar{p}}$ [see Eq. (10)] with $\Delta_{p}=0.2\hbar k_{0}$. (b) Time evolution of $O_{\bar{p}}$ [see Eq. (12)] with $\Delta^{\prime}_{p}=0.03\hbar k_{0}$. For the two curves with the highest $\bar{p}$ in (b), $\bar{p}=0.33\hbar k_{0}$ and $0.39\hbar k_{0}$, $\Delta^{\prime}_{p}=0.06\hbar k_{0}$ and we report $O_{\bar{p}}$ divided by 2. The hydrodynamization coherence time, see Eq. (9), is $T^{c}_{\rm hd}=66\mu s$.
In Ref. [5], a pulse with $t_{\rm pulse}=6$ $\mu$s and $U_{\rm pulse}=27$ $E_{r}$ was applied to systems with different values of $U_{0}$ (2.6 $E_{r}$, 6.5 $E_{r}$, and 22.9 $E_{r}$), which correspond to different values of $\bar{\gamma}_{0}$ (3.4, 2.3, and 0.94). The redistribution of energy among momentum modes was probed by studying the integrated kinetic energy per particle
[…]
within the momentum intervals $\Delta_{p}$ centered around $\bar{p}$. Figure 1(a) (adapted from Ref. [5]) shows that $E_{\bar{p}}$ exhibits rapid changes across the momentum modes on the hydrodynamization time scale. For modes near the Bragg peaks this predominantly takes the form of damped oscillations with a frequency $2\omega^{c}_{\rm hd}$, while for previously nearly unoccupied modes with momentum near ±$\pm\hbar k$ there is an overdamped oscillation in which most of the energy transfer occurs in $t_{\rm ev}\lesssim T^{c}_{\rm hd}/2$. Theoretical calculations in the TG limit show complementary behavior in position space. The $\lambda/2$-scale variations of density generated by the Bragg pulse undergo damped oscillations with a frequency $\omega^{c}_{\rm hd}$ [5]. One can understand that the $E_{\bar{p}}$ oscillation frequency is twice the spatial one because the phase of the underlying density distribution does not affect $f(p)$.
The damping of hydrodynamization oscillations in both position and momentum are a consequence of dephasing due to the spread ($\theta_{F}$) in rapidities of the Bragg peaks. For the TG gas, the bandwidth of energy differences between the $n=0$ and ±$\pm 1$ Bragg peaks associated with $\theta_{F}$ is $W=8\hbar k\theta_{F}/(2m)$, so the corresponding dephasing time scale is
$T_{\rm dp}=\frac{2\pi\hbar}{W}=\frac{\pi m}{2k\theta_{F}}.$ (11)
Different bare-particle observables are associated with different combinations of quasiparticle operators. The specific way in which an observable evolves during hydrodynamization depends on the range, amplitudes, and relative phases of quasiparticle states that contribute to that observable. Therefore, we expect the hydrodynamization oscillations to damp in a time $T^{d}_{\rm hd}\propto T_{\rm dp}$ for all observables, but with an observable-dependent coefficient. In Ref. [5], we found the damping of hydrodynamization oscillations for $E_{\bar{p}}$ to occur in similar ways in similar times in the experiments and in the TG theory. For both experiment and TG theory, the evolution is fastest for $E_{\bar{p}}$ of the originally little-occupied modes with momentum near ±$\pm\hbar k$. This suggests that those modes have relatively equal contributions from quasiparticle states across the rapidity range. With $\theta_{F}$ as large as $\hbar k$ in that work, the spread in hydrodynamization oscillation frequencies can be as large as $\omega^{c}_{\rm hd}$, producing an overdamped oscillation in which $T^{d}_{\rm hd}\simeq T^{c}_{\rm hd}$.
Local prethermalization in the TG gas, which for $f(p)$ describes the exchange of particles among nearby momentum modes, is also a consequence of dephasing due to the spread ($\theta_{F}$) in rapidities of the Bragg peaks. Therefore, we also expect the local prethermalization time $T_{\rm lp}$ to be proportional to $T_{\rm dp}$ for all observables, as it is for the damping of hydrodynamization. However, the observable-dependent details of the time evolution are in general different from those associated with the damping of hydrodynamization. The fastest possible local prethermalization time is $T_{\rm dp}$. This is the nearly momentum-independent prethermalization time of $f(p)$ that we observe at high momenta. For momentum modes well within the central peak, the prethermalization time increases as $p$ approaches 0 (see Sec. V.1).
In Ref. [5], we found the prethermalization behavior of $f(p)$ to be qualitatively similar in the experiment and in the TG gas theory. In Fig. 1(b) (adapted from Ref. [5]), we plot the occupation per particle
[…]
within the momentum intervals $\Delta^{\prime}_{p}$ centered around $\bar{p}$, measured experimentally. The experimental equilibration times of $O_{\bar{p}}$ increase as $\bar{p}$ decreases, as is also the case in the TG gas theory. It is also interesting to note that the local prethermalization times were nearly 3 times longer in the experiments, suggesting that there is a dependence on the strength of the contact interaction $g$ [see Eq. (1)] that is yet to be understood theoretically.
Summarizing our discussion in this section, we see that damping of hydrodynamization $T^{d}_{\rm hd}$ and local prethermalization $T_{\rm lp}$ both scale like $T_{\rm dp}$, and they can both be understood with reference to dephasing quasiparticle states. They are, however, physically distinct processes. We will study these differences in the rest of this paper, taking advantage of our ability to turn on and off hydrodynamization, both in experiments and theory, and our ability to change both $k$ and $\theta_{F}$ in our theoretical calculations.
III Hydrodynamization and the lack thereof in experiment
III.1 A new single Bragg pulse
In the experimental results reported in this work, we use different pulses and pulse sequences to explore the hydrodynamization process and its absence under different conditions. Our single Bragg pulse has $t_{\rm pulse}=10$ $\mu$s and $U_{\rm pulse}=18$ $E_{r}$, and it is applied to a system that is initially in a trap with $U_{0}=2.6$ $E_{r}$ and $\bar{\gamma}_{0}=3.4$. In Fig. 2(a), we plot our experimental results for the momentum distribution right after the Bragg pulse ($t_{\rm ev}=0$), and at three later times. The most salient observation from these curves is that the momentum peaks broaden with time. In Figs. 2(b) and 2(c), we show theoretical results for the rapidity and momentum distributions, respectively, following the same Bragg pulse. The theory is for a single 1D gas in the TG limit with the same trap parameters as in the experiment. We set the atom number in the TG gas to be $N=29$, so that the energy density per particle before the pulse is the same as in the experimental system for the bundle of tubes. One can see that the rapidity distribution [Fig. 2(b)] does not evolve significantly during the time in which the momentum distribution [Fig. 2(c)] does evolve. Since the experimental system has an array of tubes with different atom numbers and finite $\gamma$, we do not expect our theoretical results for a single tube with $\gamma=\infty$ (initially in the ground state) to quantitatively match the experiment. Our goal is to espouse a simple theoretical description that allows us to qualitatively describe and understand the experimentally observed phenomena.
@@ III.1 A new single Bragg pulse
Figure 2: Time evolution of the momentum and rapidity distributions after a 10 $\mu$s Bragg pulse. (a) Experimental results for the momentum distribution for $\bar{\gamma}_{0}=3.4$. (b) and (c) Corresponding theoretical results for the rapidity (b) and momentum (c) distribution of a single 1D gas in the TG limit. In the theory, the trap parameters and the average energy per atom are the same as in the experiment before the quench.
@@
In Fig. 3(a), we plot our experimental results for $E_{\bar{p}}$ [see Eq. (10)] as a function of $t_{\rm ev}$ for momentum intervals $\Delta_{p}=0.2\hbar k$. As in the results in Ref. [5] [shown in Fig. 1(a)], for which a shorter but stronger Bragg pulse was used, the $E_{\bar{p}}$ curves exhibit two clear features on the hydrodynamization coherence time scale $T^{c}_{\rm hd}=66$ $\mu$s [see Eq. (9)]. First, the curves with intermediate $\bar{p}$ between the $n=0$ and ±$\pm 1$ Bragg peaks ($\bar{p}$ from $0.5\hbar k$ to $1.5\hbar k$) rise rapidly. Second, there are oscillations with a period that is close $T^{c}_{\rm hd}/2$; these are most clearly seen in the curves with $\bar{p}$ close to 00 or $2\hbar k$. The kinetic energy curves in Fig. 3(a) are qualitatively similar to those in Fig. 1(a), illustrating that the hydrodynamization process is robust. Small changes arise mostly because the difference in the experimental parameters modify $T_{\rm dp}$ (when $\bar{\gamma}_{0}$ is greater, $\theta_{F}$ is smaller).
@@
Figure 3: Hydrodynamization after a 10 $\mu$s Bragg pulse quench. Each curve shows the time evolution of the integrated energy $E_{\bar{p}}$ over a $0.2\hbar k$ wide momentum group [see Eq. (10)]. (a) Experimental results for $\bar{\gamma}_{0}=3.4$. (b) Theory calculation for a single 1D gas in the TG limit with the same trap parameters and average energy per atom as in the experiment before the quench. The dashed lines are guides to the eye.
@@
In Fig. 3(b), we plot results from our numerical simulations of a single 1D gas in the TG limit with the same trap parameters as in the experiment. The theoretical $E_{\bar{p}}$ curves capture the main experimental features. One can see both the rapid increase in the energies of the modes near $\bar{p}\sim\hbar k$, and the oscillations in energies of the modes with $\bar{p}$ close to 00 or $2\hbar k$. Evidently, averaging over an array of 1D gases does not change the qualitative behavior of what occurs in a single tube. We note that the changes in the experimental curves are about half as large as those in the theoretical curves, possibly a result of averaging and possibly simply because the relationships between rapidities and momenta depend on the strength $g$ of the contact interaction.
III.2 QNC sequence
Equation (8) shows that each quasiparticle is in a superposition of plane waves with rapidities differing by $n2\hbar k$. It also shows that the energy distribution of the quasiparticles after the Bragg pulse is multimodal, with peaks centered at energies $(n2\hbar k)^{2}/2m$. For the rapidity distribution obtained by modeling the experiment [see Fig. 2(b)], these peaks are well separated. As we argued using Eq. (8), the energy difference between the most populated $n=0$ and ±$\pm 1$ Bragg peaks is the one that determines $T^{c}_{\rm hd}$. We therefore expect that if the quasiparticles after the quench were to be in superpositions of plane waves with similar energies, say with $n=\pm 1$, there will be no hydrodynamization. In such a situation, the state after the quench is still “high-energy” compared to the ground state, since $(2\hbar k)^{2}/2m$ remains the highest energy scale, yet there are no dynamics associated with that energy scale. The goal of this section is to confirm this expectation, showing that hydrodynamization is not an automatic outcome of dynamics after “high-energy” quenches.
A state in which the most populated Bragg peaks are the ones with $n=\pm 1$ can be realized experimentally by applying a QNC sequence [24, 13]. The QNC sequence contains two Bragg pulses like the ones in Eq. (7). Optimal QNC pulses are $\pi/(4\sqrt{2}\omega_{r})$ long (23 $\mu$s for $k_{0}$), separated by $\pi/(4\omega_{r})$ (33 $\mu$s for $k_{0}$), where $\omega_{r}=E_{r}/\hbar$ [24, 13]. These pulses are a large enough fraction of $T^{c}_{\rm hd}$ that significant hydrodynamization can occur during them. To reduce this effect, we use asymmetric QNC pulses. A long pulse followed by a short pulse still largely depletes the central peak, but at the cost of requiring a higher pulse amplitude, which compromises the two-level approximation that underlies the theory [24] and leads to a larger excitation of higher Bragg peaks. The advantage, however, is that there is less many-body evolution during the second pulse, so that the evolution after the second pulse presents a cleaner test of the absence of hydrodynamization. We label the amplitude and duration of the first (second) Bragg pulse as $U_{\rm pulse;1}$ and $t_{\rm pulse;1}$ ($U_{\rm pulse;2}$ and $t_{\rm pulse;2}$), respectively. We calculate these parameters assuming an ideal two-level system [24], which ignores many-body interactions. The ideal pulse separation is always the same as for symmetric pulses. For simplicity, we set $U_{\rm pulse;1}=U_{\rm pulse;2}$.
Figure 4: Momentum and rapidity distributions at the end of different QNC sequences. Results are shown after three sequences with: $t_{\rm pulse;1}=21$ $\mu$s, $t_{\rm mid}=33$ $\mu$s, and $t_{\rm pulse;2}=8$ $\mu$s (green dashed-dotted line); $t_{\rm pulse;1}=23$ $\mu$s, $t_{\rm mid}=33$ $\mu$s, and $t_{\rm pulse;2}=10$ $\mu$s (blue solid line); and $t_{\rm pulse;1}=24$ $\mu$s, $t_{\rm mid}=33$ $\mu$s, and $t_{\rm pulse;2}=12$ $\mu$s (red dashed line). (a) Experimental results for the momentum distribution. (b) and (c) Corresponding theoretical results for the rapidity (b) and momentum (c) distribution of a single 1D gas in the TG limit.
The length of the second pulse represents a trade-off between it being short and minimizing the $|n|=2$ excitations. Because many-body interactions complicate the ideal two-level system results in a way that is hard to calculate, we empirically determine a suitable QNC sequence. We start from the same pre-quench state (which is also the same as for the single Bragg pulse discussed in the previous subsection), and change the duration (and amplitude) of the two Bragg pulses. Figures. 4(a)–4(c) show respectively results for the momentum distributions obtained experimentally, and the rapidity and momentum distributions obtained theoretically, each after three different QNC sequences. The theoretical momentum distributions are qualitatively similar to the experimental ones, although the theoretical results exhibit lower and broader peaks than the experiment. Of the three QNC sequences, we choose to work with $t_{\rm pulse;1}=23$ $\mu$s, $t_{\rm mid}=33$ $\mu$s, and $t_{\rm pulse;2}=10$ $\mu$s (blue solid line). When $t_{\rm pulse;2}$ is shorter, the occupation of the $n=\pm 1$ Bragg peaks starts to drop in favor of higher $|n|$ peaks. Still, one can see that both in the experiment and the theory there are a nonnegligible number of atoms [and of quasiparticles in Fig. 4(b)] populating the $n=\pm 2$ Bragg peaks. We further note from Fig. 4(b) that although the QNC sequences transfer nearly all quasiparticles with $\theta\sim 0$ to higher rapidity modes, some quasiparticles with rapidities $\theta\lesssim\theta_{F}$ remain. They are the reason there are small low-momentum side peaks in the experimental and theoretical momentum distributions about $p=0$.
Figure 5: Suppression of hydrodynamization after a QNC sequence. The curves show the time evolution of $E_{\bar{p}}$, in $0.2\hbar k$ wide momentum intervals, following the QNC sequence with $t_{\rm pulse;1}=23$ $\mu$s, $t_{\rm mid}=33$ $\mu$s, and $t_{\rm pulse;2}=10$ $\mu$s. On the $x$ axis, $t_{\rm ev}=0$ is the time at which the second Bragg pulse ends. We also plot, at $t_{\rm ev}=-10$ $\mu$s, the values of $E_{\bar{p}}$ at the beginning of the second pulse (the second pulse duration is indicated by the vertical gray band). (a) Experimental results for $\bar{\gamma}_{0}=3.4$. (b) Corresponding theoretical results for a single 1D gas in the TG limit, with the same trap parameters and energy density per particle as the experimental array of 1D gases. Dashed lines are guides to the eye.
In Fig. 5, we plot the time evolution of $E_{\bar{p}}$ after the 23-33-10 $\mu$s QNC sequence in the experiment [Fig. 5(a)] and the corresponding theoretical results for a single 1D gas in the TG limit [Fig. 5(b)]. The grey shaded area indicates the time of the second ($t_{\rm pulse;2}=10$ $\mu$s) Bragg pulse in the sequence, which has the same amplitude and duration as the single Bragg pulse after which the dynamics are shown in Fig. 3. Figure 5 also shows, at $t_{\rm ev}=-10$ $\mu$s, the values of $E_{\bar{p}}$ at the beginning of the second pulse. The dominant effect of the second pulse is to deplete the $n=0$ peaks in the rapidity and momentum distributions. It also affects the occupation of the momentum modes between the $n=0$ and $|n|=1$ peaks, partially undoing the hydrodynamization evolution after the first pulse. Qualitatively, the system is far from equilibrium after the second pulse, as it would be after a single pulse of the same length. It just lacks the $n=0$ peak.
Figure 6: Hydrodynamization at higher momenta. Time evolution of $E_{\bar{p}}$ [see Eq. (10)] between $2\hbar k_{0}$ and $4\hbar k_{0}$. The QNC sequence is the same as in Fig. 5, and we use $0.2\hbar k$ wide momentum intervals to compute $E_{\bar{p}}$. (a) Experimental results for $\bar{\gamma}_{0}=3.4$. (b) Corresponding theoretical results for a single 1D gas in the TG limit, with the same trap parameters and energy density per particle as the experimental array of 1D gases. Dashed lines are guides to the eye.
The experimental and the theoretical results for $E_{\bar{p}}$ in Figs. 5(a) and 5(b), respectively, are qualitatively similar to each other for $t_{\rm ev}>0$. The curves are all relatively flat, with no oscillations with $T^{c}_{\rm hd}/2=33$ $\mu$s, in stark contrast to the results due to the single pulse in Fig. 3. When the central peak is suppressed, so is hydrodynamization.
The curves in Fig. 5 feature oscillations with a higher frequency than that in Fig. 3. Such short time scale momentum mode changes are caused by the presence of atoms with $n=\pm 2$ (see Fig. 4). In Fig. 6, we show our experimental [Fig. 6(a)] and theoretical [Fig. 6(b)] results for $E_{\bar{p}}$ of momentum modes between $2\hbar k_{0}$ and $4\hbar k_{0}$. At those energies, we find a stronger hydrodynamization signal at the frequency associated with the difference between the energies of the $n=\pm 2$ and ±$\pm 1$ Bragg peaks, which is three times higher than that between the $n=\pm 1$ and 00 Bragg peaks. The theoretical data in particular, which has better time resolution, exhibit clear oscillations of $E_{\bar{p}}$ with the expected frequency.
There is no QNC pulse sequence that allows us to explore what happens when there are only $|n|=1$ peaks. In Sec. V, we solve this problem by directly creating theoretical initial states in which the quasiparticles are in finite superpositions of plane waves with only the $n=\pm 1$ peaks present, and we study their dynamics under a translationally invariant Hamiltonian.
IV Post-quench state modeling

### v3

@@ A.1 Experimental setup
A.2 Tonks-Girardeau theory
We carry out our theoretical analyses in the strong coupling Tonks–Girardeau (TG) limit ($\gamma\to\infty$) of Hamiltonian (9), where out-of-equilibrium momentum and rapidity distributions can be readily calculated for arbitrary initial states. As shown in Ref. [8] and in Appendix B, the theoretical results in that limit parallel the experimental observations of hydrodynamization and local prethermalization at finite $\gamma$. In our calculations, we use the low-density regime of the lattice hard-core boson Hamiltonian,
${\cal H}_{\rm HCB}=-J\sum_{j=1}^{N_{s}-1}\big{(}\hat{b}^{\dagger}_{j+1}\hat{b}_{j}+{\rm H.c.}\big{)}+\sum_{j=1}^{N_{s}}U(z_{j})\hat{b}^{\dagger}_{j}\hat{b}_{j}\,,$ (11)
where $J$ is the hopping amplitude of a hard-core boson, and $N_{s}$ is the total number of lattice sites. The boson creation (annihilation) operator $\hat{b}^{\dagger}_{j}$ ($\hat{b}_{j}$) creates (annihilates) a boson at lattice site $j$, and the hard-core constraint is enforced by $\hat{b}_{j}^{2}=\hat{b}^{\dagger 2}_{j}=0$. The position of site $j$ on the lattice is set to be $z_{j}=(j-N_{s}/2)a$, with $a$ the lattice spacing. In the limit in which the site occupation $n_{j}=\langle\hat{b}^{\dagger}_{j}\hat{b}_{j}\rangle\to 0$ on all sites, the lattice hard-core boson Hamiltonian (11) is equivalent to the continuum Hamiltonian (9) in the TG limit [47]. The parameters of these two Hamiltonians are connected by $J=\hbar^{2}/(2ma^{2})$.
The main experimental observable is the momentum distribution $f(p)$, which is measured via a time-of-flight expansion [48] after the 1D interactions have been essentially turned off [15]. The redistribution of energy among momentum modes is the probed by studying the integrated kinetic energy per particle
[…]
within the momentum intervals $\Delta_{p}$ centered around $\bar{p}$. For the TG theory, $f(p)$ is written as (up to a normalization constant)
@@ A.2 Tonks-Girardeau theory
A.3 Bragg pulses
To create out-of-equilibrium states that feature hydrodynamization, in Ref. [8] and in Appendix B, we apply a Bragg pulse quench to the system with a $\lambda$ wavelength standing wave. That is, we turned on an additional potential,
@@ A.3 Bragg pulses
with wave number $k$ and amplitude $U_{\rm pulse}$, for a short period of time $t_{\rm pulse}$. In our experiments $k\approx k_{0}$, where $k_{0}$ is the wave number of the 2D lattice. After the pulse, the system was allowed to evolve for different times $t_{\rm ev}$ before the momentum distribution $f(p)$ was measured.
The effect of the Bragg pulse can be understood analytically in the Raman-Nath regime ($t_{\rm pulse}\to 0$) for the homogeneous TG gas. The pulse in this regime, known as a Kapitza-Dirac pulse, “kicks” the particles and quasiparticles before they have had time to move. Such a kick generates a state in which the quasiparticles are in a superposition of plane waves that differ by even integer multiples of the Bragg momentum [51]. The many-quasiparticle (noninteracting fermion) wave function changes from that of a Fermi sea $|\Psi_{0}\rangle=\prod_{|\theta|<\theta_{F}}\hat{c}^{\dagger}_{\theta}|0\rangle$ to
@@
$I_{n}(-ib)$ are the modified Bessel functions of the first kind, where the integer $n$ is the order of the Bragg peak and $b$ depends on the area of the Bragg pulse. In a homogeneous system, the rapidity distribution $f(\theta)$ is conserved after the quench, but the momentum distribution $f(p)$ for the bare particles evolves in time. The time evolution of the relative phases of different $\hat{c}^{\dagger}_{\theta}$ enters $f(p)$ via the Bose-Fermi mapping, so $f(p)$ changes in time despite $f(\theta)$ being constant. That is, although there is no redistribution of energy among the quasiparticles, there is a redistribution of energy among different momentum modes of the particles.
Appendix B A new single Bragg pulse
In the experimental results reported in this work, we use different pulses and pulse sequences to explore the hydrodynamization process and its absence under different conditions. Our single Bragg pulse has $t_{\rm pulse}=10$ $\mu$s and $U_{\rm pulse}=18$ $E_{r}$, and it is applied to a system that is initially in a trap with $U_{0}=2.6$ $E_{r}$ and $\bar{\gamma}_{0}=3.4$. We choose these Bragg pulse parameters to correspond to the second Bragg pulse of the Newton’s cradle sequence described in Sec. V, which allows us to most directly compare the evolution of two- and three-peak experimental quenches.
@@ Appendix B A new single Bragg pulse
Figure 12: Time evolution of the momentum and rapidity distributions after a 10 $\mu$s Bragg pulse. (a) Experimental results for the momentum distribution for $\bar{\gamma}_{0}=3.4$. (b), (c) Corresponding theoretical results for the rapidity (b) and momentum (c) distribution of a single 1D gas in the TG limit. In the theory, the trap parameters and the average energy per atom are the same as in the experiment before the quench.
In Fig. 12(a), we plot our experimental results for the momentum distribution right after the Bragg pulse ($t_{\rm ev}=0$), and at two later times. The most salient observation from these curves is that the momentum peaks broaden with time. In Figs. 12(b) and 12(c), we show theoretical results for the rapidity and momentum distributions, respectively, following the same Bragg pulse. The theory is for a single 1D gas in the TG limit with the same trap parameters as in the experiment. We set the atom number to be $N=29$, so that the energy per particle in the ground state of the trapped TG gas before the pulse is the same as in the experimental system for the bundle of tubes. The rapidity distribution [Fig. 12(b)] does not evolve significantly during the time in which the momentum distribution [Fig. 12(c)] evolves. Since the experimental system has an array of tubes with different atom numbers and finite $\gamma$, we do not expect our theoretical results for a single tube with $\gamma=\infty$ (initially in the ground state) to quantitatively match the experiment [16]. Our goal is to espouse a simple theoretical description that allows us to qualitatively describe and understand the experimentally observed phenomena.
@@
In Fig. 13(a), we plot our experimental results for $E_{\bar{p}}$ [see Eq. (12)] as a function of $t_{\rm ev}$ for momentum intervals $\Delta_{p}=0.2\hbar k_{0}$. As in the results in Ref. [8], for which a shorter but stronger Bragg pulse was used, the $E_{\bar{p}}$ curves exhibit two clear features on the hydrodynamization coherence timescale $T^{c}_{\rm hd}=66$ $\mu$s [see Eq. (2)]. First, the curves with intermediate $\bar{p}$ between the $n=0$ and ±$\pm 1$ Bragg peaks ($\bar{p}$ from $0.5\hbar k_{0}$ to $1.5\hbar k_{0}$) rise rapidly. Second, there are oscillations with a period that is close $T^{c}_{\rm hd}/2$; these are most clearly seen in the curves with $\bar{p}$ close to 00 or $2\hbar k_{0}$. The kinetic energy curves in Fig. 13(a) are qualitatively similar to those in Ref. [8], illustrating that the hydrodynamization process is robust. Small changes arise mostly because the difference in the experimental parameters modify $T_{\rm dp}$ (when $\bar{\gamma}_{0}$ is greater, $\theta_{F}$ is smaller).
@@
Figure 13: Hydrodynamization after a 10 $\mu$s Bragg pulse quench. Each curve shows the time evolution of the integrated energy $E_{\bar{p}}$ over a $0.2\hbar k_{0}$ wide momentum group [see Eq. (12)]. (a) Experimental results for $\bar{\gamma}_{0}=3.4$. (b) Theory calculation for a single 1D gas in the TG limit with the same trap parameters and average energy per atom as in the experiment before the quench. The dashed lines are guides to the eye.
@@
In Fig. 13(b), we plot results from our numerical simulations of a single 1D gas in the TG limit with the same trap parameters as in the experiment. The theoretical $E_{\bar{p}}$ curves capture the main experimental features. One can see both the rapid increase in the energies of the modes near $\bar{p}\sim\hbar k_{0}$, and the oscillations in energies of the modes with $\bar{p}$ close to 00 or $2\hbar k_{0}$. Evidently, averaging over an array of 1D gases does not change the qualitative behavior of what occurs in a single tube. We note that the changes in the experimental curves are about half as large as those in the theoretical curves, possibly a result of averaging and possibly simply because the relationships between rapidities and momenta depend on the strength $g$ of the contact interaction.
Appendix C Insensitivity to the trap and
the Bragg pulse